Why Equillibrium constant is said to be independent of the directionfrom which equillibrium is approached? When we know that the constant gets inversed when the direction is reversed.

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The Equilibrium Constant, Kc, relates to a chemical reaction at equilibrium. It can be calculated if the equilibrium concentration of each reactant and product in a reaction at equilibrium is known. The equilibrium expression below, formed from the general chemical equilibrium, is universally true. The chemical components happen to be gases.

aA(g) + bB(g) = cC(g) + dD(g)

The Ideal Gas Equation shows that the pressure of a gas is proprtional to its concentration.

pV = nRT

where p is pressure of a particular gas (its partial pressure) in an equilibrium mixture, V is the total volume, n is the number of moles of the particular gas, R is the general gas constant, and T is the absolute temperature.

In the above equation where the temperature is also constant,

P a n/V

The equilibrium constant can therefore also be expressed in terms of partial pressures, and is denoted as Kp:

There are several types of equilibrium constants. Each is constant at a constant temperature. For example, consider the following ionic equilibrium involving the aqueous solution of a weak acid:

CH3COOH(aq) + H2O(l) = CH3COO-(aq) + H3O+(aq)

Ka is called the 'acid dissociation constant'. Note that water is omitted from the expression because it is present in such vast excess that its concentration changes negligibly on the formation of equilibrium and is therefore effectively constant. The concentration of the water is included in the equilibrium constant, and Ka can be thought of as a modified equilibrium constant.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com